The volume of 300 K sample of gas is raised from 3.0 L to 4.5 L. What will be the new temperature of the gas?


1
Expert's answer
2021-04-29T07:21:32-0400

"\frac {V_1}{T_1} =\frac {V_2} {T_2}"


"T_2=\frac {300K×4.5 L} {3.0L}"


"T_2=450K"